OpenStudy (anonymous):

V=square root of 5/2 times 6. Find v. Answer in standard radical form. I really don't get this, please help and explain.

OpenStudy (theopenstudyowl):

V=sqrt(5/2)(6)=????? V^2=5/2 times 36 V^2=90 V= sqrt 90 V= 3 sqrt 10 or 3 radical 10

OpenStudy (theopenstudyowl):

make sense?

OpenStudy (anonymous):

Wait where did 36 came from?

OpenStudy (theopenstudyowl):

because we squared "V".

OpenStudy (theopenstudyowl):

radical 5/2 squared is 5/2 and 6 squared is 36... ,ake sense?

OpenStudy (anonymous):

Why did we square v?

OpenStudy (theopenstudyowl):

ok, I am going to write a solution using the equation system on openstudy, can u gie me like 5 minutes?

OpenStudy (anonymous):

Okay

OpenStudy (theopenstudyowl):

\[V=\sqrt{5/2}*6, V^2=(\sqrt{5/2})^2*(6)^2, V^2=5/2*36, V^2=90,V=\sqrt{90}, V=(\sqrt{9}*\sqrt{10}, V=3\sqrt{10}\]

OpenStudy (theopenstudyowl):

pellet it didnt work right

OpenStudy (theopenstudyowl):

well tell me if it makes sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is that it?

OpenStudy (theopenstudyowl):

No, not everything showed up Sorry only part of the solution

OpenStudy (anonymous):

Okay so what's next?